Our body is a moderately good conductor of electricity. Taken as a whole, our body’s resistance measured for conduction from the fingers of one hand to the other is about 1500 ohms when dry and 500 ohms when wet. The electric potential at a wall socket is 220 V, so if we inadvertently touch a high potential house with one hand, what will be the current trough our body when (a) dry and (b) wet. (c) Using the chart below, what do you think will happen to us if there are such amounts of current passing through our body?


1
Expert's answer
2021-03-18T15:24:03-0400

Given,

Resistance of our body when dry"(R_{dry})=1500\Omega"

Resistance of our body when wet "(R_{wet})=500\Omega"

Electrical potential at a wall socket "(V)=220V"

a) Current in the body when body is dry,

"(I)=\frac{V}{R_{dry}}"


"(I)=\frac{220}{1500}A"


"(I)=0.146A"


b) When body is wet, then current "(I)=\frac{V}{R_{wet}}"

"\Rightarrow I=\frac{220}{500}A"

"\Rightarrow I =0.44A"

c) When electric current passes through our body, then

  • It can be able to create severe burn in our body.
  • Because of electric shock, our body may face cardiac arrest.
  • Burn to tissues and organs.
  • muscle spasms.
  • serious effects to the nervous system and other unexpected consequences